# Weighted Parameter Estimators of the Generalized Extreme Value Distribution in the Presence of Missing Observations

### v1

@@ 1 Introduction
Various techniques have been developed to analyze extreme values in the presence of missing observations. In Beirlant et al. (2023); Xu et al. (2022); Zou et al. (2020), authors examined the Hill estimator under various missing data and distributional assumptions. Specifically, these works considered the missing data scenario where the largest order statistics in the observed data set were missing. In contrast, Ayiah-Mensah et al. (2021) and Turki et al. (2020) both considered different types of multiple imputation to fill in the missing observations in their extreme value analysis. This multiple imputation approach implicitly assumes the probability of an observation being missing is completely independent of the observed data or depends only on the observed data. In Rydén (2008), the author examined the missing data problem from an applied perspective by studying return values of wave height whereas in Glavas̆ and Mladenović (2024), the authors examined the asymptotic properties of moving averages from the Gumbel distribution in the presence of missing observations. Alternatively, the analysis of Beck et al. (2020) simply removed the blocks of observations with high levels of missingness according to some preset threshold and treated all other maxima as though they were derived from fully observed blocks of data. While this approach seems to be used more often in practice, there is no general consensus on an optimal technique for handling missing records.
@@ 2 Block Maxima Approach to Extremes
where $-\infty<\mu<\infty$, $\sigma>0$, and $-\infty<\xi<\infty$ Fisher and Tippett (1928); Gnedenko (1943); Von Mises (1936). Note that this result also holds for certain stationary series, specifically those satisfying the condition that the form of the temporal dependence results in extreme events that are approximately independent if they are sufficiently far apart. See Leadbetter (1974) for relevant theoretical results and Beirlant et al. (2004) for an in-depth explanation of the extension of this result to stationary series, and estimation.
In practice, we approximate the distribution of $M_{n}$ by $G$ for large enough $n$. The parameters of the GEV distribution are typically estimated by one of two methods: (i) maximum likelihood (ML) estimation or (ii) probability weighted moment estimators (PWM). For sake of simplicity in the estimation descriptions below, we suppose the maxima $M_{1}$, …, $M_{k}$ are independently and identically distributed according to a GEV distribution.
@@ 2.1 Maximum Likelihood Estimation
$\hat{\Theta}=\operatorname{argmax}_{\Theta}\ell(\Theta;m_{1},...,m_{k})=\operatorname{argmax}_{\Theta}\sum_{j=1}^{k}\log\left\{g(m_{j};\Theta)\right\},$ (3)
@@ 3.1 Missingness in the block maxima framework
First, we slightly modify the previous notation to account for missing observations. Suppose of the series of $N$ random variables, $N_{obs}$ are observed and $N_{miss}$ are missing (i.e. $N_{obs}+N_{miss}=N$). Let $\mathscr{S}$ and $\mathscr{S}^{\prime}$ denote the disjoint sets of indices of sizes $N_{obs}$ and $N_{miss}$ for the observed and missing random variables, respectively, such that $\mathscr{S}_{N}=\mathscr{S}\cup\mathscr{S}^{\prime}=\{1,\dots,N\}$. We further let $\{X_{i}:i\in\mathscr{S}\}$ and $\{X_{i}^{\prime}:i\in\mathscr{S}^{\prime}\}$ denote the corresponding sets of observed and missing random variables all with CDF $F(\cdot)$. Again, we suppose $N$ is partitioned into $k$ blocks each of size $n$, $kn=N$. In the $j$th block, denote the maxima based on the $n_{j}$ observed and $n_{j}^{\prime}$ missing observations as $M_{n_{j}}$ and $M_{n_{j}^{\prime}}$, respectively, and true block maxima as $M_{j}$, where $n_{j}+n_{j}^{\prime}=n$.
@@
Using the missingness notation within the block maxima framework, the three categories of missingness can be described probabilistically as follows. Suppose $I$ is an indicator of missingness for a random variable $X$, where $I=1$ if $X$ is observed and 0 otherwise.
@@ 1. MCAR:
    The probability of missingness is fixed across all observation in the series, i.e., $P(I=0\mid\{X_{i}:i\in\mathscr{S}_{N}\})=P(I=0)$. In this case, each block may or may not have a portion of missing observations.
@@ 3. MNAR:
In practice, it is more likely to encounter missingness mechanisms that fall into the MAR and MNAR categories, rather than MCAR. In the context of wave surges, missingness being completely independent of wave intensity or any other information from the observed/unobserved data would indicate MCAR. In contrast, increased failure of the measuring equipment over time would be an example of MAR, and failure of equipment to measure high intensity surges would be an example of MNAR. In Figure 1, we illustrate how different missingness mechanisms affect the observed block maxima data. Specifically, we generated 75 draws from an Exponential distribution with rate parameter $0.2$ and applied either an MCAR, MAR or MNAR missingness mechanism in the left, middle and right panels, respectively. In the left most panels, we assumed an MCAR missingness mechanism where series level data were missing according to independent Bernoulli draws with probability equal to $0.3$. Under this missingness mechanism, a particular block’s maximum will be missing with probability $0.3$. In the lower left panel of Figure 1, blocks 1 to 3 and 5 still contain their true maxima whereas block 4’s observed maxima now corresponds to the second largest maxima in the block. Similarly, in the middle panels, we imposed a MAR mechanism where the series level are missing with increasing probability according to their time indices (i.e. an observation’s probability of missing increases according to a Normal cumulative distribution function). Thus, the earlier blocks have fewer missing observations whereas later blocks have mostly missing observations. The corresponding effect on an extreme value analysis is that the observed maxima in earlier blocks are more likely to correspond to the true block maxima whereas later blocks’ observed maxima will correspond to the lower order statistics. In contrast, in the right most panels of Figure 1, we consider the most extreme case of a MNAR mechanism where 25% of the upper order statistics in the series level data are missing. This scenario is the most destructive to an extreme value analysis as the upper order statistics on which an analysis is based are totally unobserved. In this example, all five blocks’ observed maxima correspond to much lower order statistics and will therefore yield biased inferences. We note that in any applied analysis, certain assumptions are required on the type of missingness mechanism. Thus, the described missingness mechanisms, while not necessarily applicable to all real world data scenarios, will be considered in an extensive simulation study in Section 4 as possible cases on a spectrum of how negatively missing series level data can impact block maxima inferences.
@@
We denote $\hat{w}_{j}^{(1)}$ as the unconditional estimate of not missing the true block maximum within block $j$, i.e., $\hat{w}_{j}^{(1)}=P\{M_{j}=M_{n_{j}}\}$, as it depends only on the number of observations (observed/missing) within block $j$ and not on block $j$’s data. In contrast, we denote $\hat{w}_{j}^{(2)}$ as the conditional estimate of not missing the true block maximum within block $j$, i.e. $\hat{w}_{j}^{(2)}=P(M_{j}=M_{n_{j}}\mid M_{n_{j}}=m_{n_{j}},\{X_{i}\}_{i\in\mathscr{S}})$, as it depends on all non-missing observations across all blocks. Throughout the remainder of the article, we will assume that any analyses are conducted conditionally on the number of missing and non-missing observations within each of the blocks. Additional commentary regarding unconditional analyses is included in Section 6.
We note that although these weighting schemes are inspired by MCAR and MAR missingness mechanisms, respectively, their use is not restricted to such cases as these missingness mechanisms are not required as assumptions for deriving the properties of corresponding estimators outlined in Section 3.4. As such, the definitions of the unconditional/conditional weights were proposed in order to account for the three missingness categories. We assess the performance of the proposed weighted estimation procedures when applied to simulated data under three missingness mechanisms in 4.
@@ 3.2 Weighted Maximum Likelihood Estimation
Weighted ML estimation is a technique of downweighting the effect of outlying observations in an extreme value analysis (see Carreau and Girard (2011), Dupuis and Morgenthaler (2002b) and Wang et al. (2010) for further details) as well as accounting for missingness through inverse probability weighting Seaman and White (2013). The weighted log-likelihood of the GEV model is defined for an appropriate choice of weights $\hat{w}_{j}$ for $j=1,\dots,k$ as follows:
@@ 3.4 Statistical properties of estimators
For the weighted likelihood estimators, when the weights are estimated unconditionally, the weighted maximum likelihood estimator for the GEV parameters will be consistent and asymptotically Normally distributed. As the log-likelihood in Equation (9) can be regarded as the relevance weighted log likelihood function where the weights correspond to the importance of the observations $m_{n_{j}}$, then when $\hat{w}_{j}^{(1)}=\frac{n_{j}}{n_{j}+n_{j}^{\prime}}$, the asymptotic arguments of Hu can be applied to $\hat{\Theta}=\operatorname{argmax}_{\Theta}\ell_{\hat{\boldsymbol{w}}}(\Theta;m_{n_{1}},...,m_{n_{k}})$ for $\xi>-0.5$ Hu (1997). In the case when the weights are estimated conditionally and are evaluated at the observed maxima, $\hat{w}_{j}^{(2)}=P(M_{j}=M_{n_{j}}\mid M_{n_{j}}=m_{n_{j}},\{X_{i}\}_{i\in\mathscr{S}})$, the asymptotic arguments of Hu (1997) do not immediately follow. An analogous data-dependent weighting approach was recently considered by in Biswas et al. (2015) however the corresponding asymptotic properties have yet to be determined. A full description on how the weighted maximum likelihood estimators, using unconditional weights, satisfy the regularity conditions stated by Hu for consistency and asymptotic normality is given in Appendix A.
@@ 4 Simulation Study
We replicated the simulation procedure over $1000$ iterations and computed the unnormalized Cramér–von Mises (CvM) distances, where the CvM distance is defined by:
@@
where $x_{i}$ corresponds to the $N^{\ast}$ incremental percentiles of $G(\cdot)$ from $2\%$ to $98\%$, which yields similar results to a Monte Carlo-based estimate.
@@
We report the t-distribution simulation results in Tables 1-3, and the Exponential and Beta distribution simulation results in Tables 1-6 in the Supplementary Materials. From Tables 1-3, the conditional-weighted ML estimation outperforms the unweighted and unconditionally weighted ML methods in all scenarios, i.e., the conditional ML estimator yielded the lowest CvM distances irrespective of whether the data were MNAR, MCAR or MAR. However, in cases of high levels of missingness, e.g., when the percentage blocks with missing observations is 80%, the conditional ML estimator tended to perform poorly. In the case of the t distribution, we found no changes in the reported simulation results when the degrees of freedom the t distribution increased from $5$ to $15$ to $40$ (i.e. as the distribution tails become lighter). The superiority of the conditional ML estimator compared to the observed and unconditional ML estimators is most likely attributable to the improved accuracy of the weight in modelling the missingness probability. Specifically, since the conditional weights are based on the series level data, the empirical CDF will yield a consistent and unbiased estimator for whether the true block maxima are missing. As the unconditional weights are based solely on the relative sizes of the observed to unobserved data within the blocks, they do not utilize any information from the series level data nor the observed maxima. The results from the PWM-based estimation are less consistent. Although the weighted PWM-based estimation generally outperforms the unweighted PWM-based estimation in the MNAR setting, there is no clear improvement in performance between the unconditional and conditional weights. In the MAR and MCAR settings, this pattern is reversed with the unweighted PWM-based estimator outperforming the weighted PWM-based estimators. Furthermore, the conditional weighted likelihood method outperforms the PWM-based estimation methods in most cases irrespective of the data generating distribution or the missingness mechanism for the simulated data. We conjecture that the weighted moment estimators’ poor behaviour is due to the reweighting of the observed maxima rather than the likelihood contributions as in the ML estimation procedure. By reweighting the observed maxima, the weighted PWM-based estimators are attempting to correct for the cases where the observed block maxima do/do not correspond to the true block maxima values. Prediction of new extreme order statistics based on the observed order statistics is highly variable and thus can lead to innaccurate predictions. Based on the stability of the ML estimator simulation results given above relative to the PWM estimator results, we would advise using the conditionally weighted ML estimator to obtain estimates of the GEV parameters series level observations are missing.
@@
Table 1: Average empirical Cramér–von Mises distances (non-normalized) between the CDFs of the estimated complete data GEV and the proposed modelling approach computed over $1000$ replications (“mle-obs” - observed likelihood, “mle-uncond” - unconditional weighted likelihood, “mle-cond” - conditional weighted likelihood”, “mom-obs” - observed moments, “mom-uncond” - unconditional weighted moments, “mom-cond” - conditional weighted moments) for MNAR data generated from an t distribution (df=5) with blocks containing $100$ observations. Simulation parameters: “sims” - number of blocks, “pbm” - proportion of blocks with missing observations, “pm” - expected proportion of missing observations within blocks.
@@
Table 2: Average empirical Cramér–von Mises distances (non-normalized) between the CDFs of the estimated complete data GEV and the proposed modelling approach computed over $1000$ replications (“mle-obs” - observed likelihood, “mle-uncond” - unconditional weighted likelihood, “mle-cond” - conditional weighted likelihood”, “mom-obs” - observed moments, “mom-uncond” - unconditional weighted moments, “mom-cond” - conditional weighted moments) for MCAR data generated from an t distribution (df=5) with blocks containing $100$ observations. Simulation parameters: “sims” - number of blocks, “pbm” - proportion of blocks with missing observations, “pm” - expected proportion of missing observations within blocks.
@@
Table 3: Average empirical Cramér–von Mises distances (non-normalized) between the CDFs of the estimated complete data GEV and the proposed modelling approach computed over $1000$ replications (“mle-obs” - observed likelihood, “mle-uncond” - unconditional weighted likelihood, “mle-cond” - conditional weighted likelihood”, “mom-obs” - observed moments, “mom-uncond” - unconditional weighted moments, “mom-cond” - conditional weighted moments) for MAR data generated from an t distribution (df=5) with blocks containing $50$ or $100$ observations. Simulation parameters: “sims” - number of blocks, “apm” - average proportion of missingness across the series.

### v2

@@ 1 Introduction
Various techniques have been developed to analyze extreme values in the presence of missing observations. In Beirlant et al. (2023); Xu et al. (2022); Zou et al. (2020), authors examined the Hill estimator under various missing data and distributional assumptions. Specifically, these works considered the missing data scenario where the largest order statistics in the observed data set were missing. In contrast, Ayiah-Mensah et al. (2021) and Turki et al. (2020) both considered different types of multiple imputation to fill in the missing observations in their extreme value analysis. This multiple imputation approach implicitly assumes the probability of an observation being missing is completely independent of the observed data or depends only on the observed data In Rydén (2008), the author examined the missing data problem from an applied perspective by studying return values of wave height whereas in Glavas̆ and Mladenović (2024), the authors examined the asymptotic properties of moving averages from the Gumbel distribution in the presence of missing observations. Alternatively, the analysis of Beck et al. (2020) simply removed the blocks of observations with high levels of missingness according to some preset threshold and treated all other maxima as though they were derived from fully observed blocks of data. While this approach seems to be used more often in practice, there is no general consensus on an optimal technique for handling missing records.
@@ 2 Block Maxima Approach to Extremes
where $-\infty<\mu<\infty$, $\sigma>0$, and $-\infty<\xi<\infty$ Fisher and Tippett (1928); Gnedenko (1943); Von Mises (1936). Note that this result also holds for certain stationary series, specifically those satisfying the condition that the form of the temporal dependence results in extreme events that are approximately independent if they are sufficiently far apart. See Leadbetter (1974) for relevant theoretical results and Beirlant et al. (2004) for an in-depth explanation of the extension of this result to stationary series and estimation.
In practice, we approximate the distribution of $M_{n}$ by $G$ for large enough $n$. The parameters of the GEV distribution are typically estimated by one of two methods: (i) maximum likelihood (ML) estimation or (ii) probability weighted moment estimators (PWM). For the sake of simplicity in the estimation descriptions below, we suppose the maxima $M_{1}$, …, $M_{k}$ are independently and identically distributed according to a GEV distribution.
@@ 2.1 Maximum Likelihood Estimation
$\hat{\Theta}={\arg\max}_{\Theta}\ell(\Theta;m_{1},...,m_{k})={\arg\max}_{\Theta}\sum_{j=1}^{k}\log\left\{g(m_{j};\Theta)\right\},$ (3)
@@ 3.1 Missingness in the block maxima framework
First, we slightly modify the previous notation to account for missing observations. Suppose of the series of independent and identically distributed $N$ random variables, $N_{obs}$ are observed and $N_{miss}$ are missing (i.e. $N_{obs}+N_{miss}=N$). Let $\mathscr{S}$ and $\mathscr{S}^{\prime}$ denote the disjoint sets of indices of sizes $N_{obs}$ and $N_{miss}$ for the observed and missing random variables, respectively, such that $\mathscr{S}_{N}=\mathscr{S}\cup\mathscr{S}^{\prime}=\{1,\dots,N\}$. We further let $\{X_{i}:i\in\mathscr{S}\}$ and $\{X_{i}^{\prime}:i\in\mathscr{S}^{\prime}\}$ denote the corresponding sets of observed and missing random variables all with CDF $F(\cdot)$. Again, we suppose $N$ is partitioned into $k$ blocks each of size $n$, $kn=N$. In the $j$th block, denote the maximum based on the $n_{j}$ observed and $n_{j}^{\prime}$ missing observations as $M_{n_{j}}$ and $M_{n_{j}^{\prime}}$, respectively, and the true block maximum as $M_{j}$, where $n_{j}+n_{j}^{\prime}=n$.
@@
Using the missingness notation within the block maxima framework, the three categories of missingness can be described probabilistically as follows. Suppose $I$ is an indicator of missingness for an arbitrary random variable $X$, where $I=1$ if $X$ is observed and 0 otherwise.
@@ 1. MCAR:
    The probability of missingness is fixed across all observations in the series, i.e., $P(I=0\mid\{X_{i}:i\in\mathscr{S}_{N}\})=P(I=0)$. In this case, each block may or may not have a portion of missing observations.
@@ 3. MNAR:
In practice, it is more likely to encounter missingness mechanisms that fall into the MAR and MNAR categories, rather than MCAR. In the context of wave surges, missingness being completely independent of wave intensity or any other information from the observed/unobserved data would indicate MCAR. In contrast, increased failure of the measuring equipment over time would be an example of MAR, and failure of equipment to measure high intensity surges would be an example of MNAR. In Figure 1, we illustrate how different missingness mechanisms affect the observed block maxima data. Specifically, we generated 75 draws from an Exponential distribution with rate parameter 0.20.20.2 and applied either an MCAR, MAR or MNAR missingness mechanism in the left, middle and right panels, respectively. In the left most panels, we assumed an MCAR missingness mechanism where series level data were missing according to independent Bernoulli draws with probability equal to 0.30.30.3. Under this missingness mechanism, a particular block’s maximum will be missing with probability 0.30.30.3. In the lower left panel of Figure 1, blocks 1 to 3 and 5 still contain their true maxima whereas block 4’s observed maxima now corresponds to the second largest maxima in the block. Similarly, in the middle panels, we imposed a MAR mechanism where the series level are missing with increasing probability according to their time indices (i.e. an observation’s probability of being missing increases according to a Normal cumulative distribution function). Thus, the earlier blocks have fewer missing observations whereas later blocks have mostly missing observations. The corresponding effect on an extreme value analysis is that the observed maxima in earlier blocks are more likely to correspond to the true block maxima whereas later blocks’ observed maxima will correspond to the lower order statistics. In contrast, in the right most panels of Figure 1, we consider the most extreme case of a MNAR mechanism where the highest 25% of the order statistics of the entire series are removed. This scenario is the most destructive to an extreme value analysis as the upper order statistics on which an analysis is based are totally unobserved. In this example, all five blocks’ observed maxima correspond to much lower order statistics and will therefore yield biased inferences. We note that in any applied analysis, certain assumptions are required on the type of missingness mechanism. Thus, the described missingness mechanisms, while not necessarily applicable to all real world data scenarios, will be considered in an extensive simulation study in Section 4 as possible cases on a spectrum of how negatively missing series level data can impact block maxima inferences.
@@
We denote $\hat{w}_{j}^{(1)}$ as the unconditional estimate of not missing the true block maximum within block $j$, i.e., $\hat{w}_{j}^{(1)}=\hat{P}\{M_{j}=M_{n_{j}}\}$, as it depends only on the number of observations (observed/missing) within block $j$ and not on block $j$’s data. In contrast, we denote $\hat{w}_{j}^{(2)}$ as the conditional estimate of not missing the true block maximum within block $j$, i.e. $\hat{w}_{j}^{(2)}=\hat{P}(M_{j}=M_{n_{j}}\mid M_{n_{j}}=m_{n_{j}},\{X_{i}\}_{i\in\mathscr{S}})$, as it depends on all non-missing observations across all blocks. Throughout the remainder of the article, we will assume that any analyses are conducted conditionally on the number of missing and non-missing observations within each of the blocks. Additional commentary regarding unconditional analyses is included in Section 6.
We note that although these weighting schemes are inspired by MCAR and MAR missingness mechanisms, respectively, their use is not restricted to such cases as these missingness mechanisms are not required as assumptions for deriving the properties of the corresponding estimators outlined in Section 3.4. As such, the definitions of the unconditional/conditional weights were proposed in order to account for the three missingness categories. We assess the performance of the proposed weighted estimation procedures when applied to simulated data under three missingness mechanisms in 4.
@@ 3.2 Weighted Maximum Likelihood Estimation
Weighted ML estimation is a technique of downweighting the effect of outlying observations in an extreme value analysis (see Carreau and Girard (2011); Dupuis and Morgenthaler (2002b); Wang et al. (2010) for further details) as well as accounting for missingness through inverse probability weighting Seaman and White (2013). The weighted log-likelihood of the GEV model is defined for an appropriate choice of weights $\hat{w}_{j}$ for $j=1,\dots,k$ as follows:
@@ 3.4 Statistical properties of estimators
For the weighted likelihood estimators, it is not immediately clear whether the weighted ML estimator for the GEV parameters will be consistent and asymptotically Normally distributed. The log-likelihood in Equation (9) can be regarded as a relevance weighted log-likelihood function as the observed block maxima are not necessarily drawn from the GEV distribution Hu (1997). The associated weights correspond to the relevance of the observed block maxima densities relative to the true block maxima density (i.e. the GEV distribution). There are multiple challenges associated in establishing the asymptotic properties of the maximum relevance weighted likelihood estimator. Depending on the assumptions underlying the missingness mechanism, it is not evident how to define the asymptotic distribution of the observed maxima as the number of observations within the block tends to infinity. Specifically, the relative position of the observed block maxima (i.e. the order of the block’s largest observed order statistic) will be dependent on the growth rate of the number of observed/unobserved random variables within the block. Furthermore, when the weights are estimated conditionally and evaluated at the observed maxima, $\hat{w}_{j}^{(2)}=\hat{P}(M_{j}=M_{n_{j}}\mid M_{n_{j}}=m_{n_{j}},\{X_{i}\}_{i\in\mathscr{S}})$, the asymptotic arguments of Hu (1997) do not immediately follow as the weights are random. An analogous data-dependent weighting approach was recently considered by Biswas et al. (2015) however the corresponding asymptotic properties have yet to be determined. A mathematical description highlighting some of the challenges stated above and the required regularity conditions stated by Hu (1997) for consistency and asymptotic normality is given in Appendix A.
@@ 4 Simulation Study
We replicated the simulation procedure over 100010001000 iterations and computed the unnormalized Cramér–von Mises (CvM) distances, where the CvM distance is defined by:
@@
[…]
@@
We report the t-distribution simulation results in Tables 1-3, and the Exponential and Beta distribution simulation results in Tables 1-6 in the Supplementary Materials. From Tables 1-3, the conditional-weighted ML estimation outperforms the unweighted and unconditionally weighted ML methods in all scenarios, i.e., the conditional ML estimator yielded the lowest CvM distances irrespective of whether the data were MNAR, MCAR or MAR. However, in cases of high levels of missingness, e.g., when the percentage blocks with missing observations is 80%, the conditional ML estimator tended to perform poorly. In the case of the t distribution, we found no changes in the reported simulation results when the degrees of freedom of the t distribution increased from 555 to 151515 to 404040 (i.e. as the distribution tails become lighter). The superiority of the conditional ML estimator compared to the observed and unconditional ML estimators is most likely attributable to the improved accuracy of the weight in modelling the missingness probability. Specifically, since the conditional weights are based on the series level data, the empirical CDF will yield a consistent and unbiased estimator for whether the true block maxima are missing. As the unconditional weights are based solely on the relative sizes of the observed to unobserved data within the blocks, they do not utilize any information from the series level data nor the observed maxima. The results from the PWM-based estimation are less consistent. Although the weighted PWM-based estimation generally outperforms the unweighted PWM-based estimation in the MNAR setting, there is no clear improvement in performance between the unconditional and conditional weights. In the MAR and MCAR settings, this pattern is reversed with the unweighted PWM-based estimator outperforming the weighted PWM-based estimators. Furthermore, the conditional weighted likelihood method outperforms the PWM-based estimation methods in most cases irrespective of the data generating distribution or the missingness mechanism for the simulated data. We conjecture that the weighted moment estimators’ poor behaviour is due to the reweighting of the observed maxima rather than the likelihood contributions as in the ML estimation procedure. By reweighting the observed maxima, the weighted PWM-based estimators are attempting to correct for the cases where the observed block maxima do/do not correspond to the true block maxima values. Prediction of new extreme order statistics based on the observed order statistics is highly variable and thus can lead to innaccurate predictions. Based on the stability of the ML estimator simulation results given above relative to the PWM estimator results, we would advise using the conditionally weighted ML estimator to obtain estimates of the GEV parameters when series level observations are missing.
@@
Table 1: Average empirical Cramér–von Mises distances (non-normalized) between the CDFs of the estimated complete data GEV and the proposed modelling approach computed over 100010001000 replications (“mle-obs” - observed likelihood, “mle-uncond” - unconditional weighted likelihood, “mle-cond” - conditional weighted likelihood”, “mom-obs” - observed moments, “mom-uncond” - unconditional weighted moments, “mom-cond” - conditional weighted moments) for MNAR data generated from an t distribution (df=5) with blocks containing 100100100 observations. Simulation parameters: “sims” - number of blocks, “pbm” - proportion of blocks with missing observations, “pm” - expected proportion of missing observations within blocks.
@@
Table 2: Average empirical Cramér–von Mises distances (non-normalized) between the CDFs of the estimated complete data GEV and the proposed modelling approach computed over 100010001000 replications (“mle-obs” - observed likelihood, “mle-uncond” - unconditional weighted likelihood, “mle-cond” - conditional weighted likelihood”, “mom-obs” - observed moments, “mom-uncond” - unconditional weighted moments, “mom-cond” - conditional weighted moments) for MCAR data generated from an t distribution (df=5) with blocks containing 100100100 observations. Simulation parameters: “sims” - number of blocks, “pbm” - proportion of blocks with missing observations, “pm” - expected proportion of missing observations within blocks.
@@
Table 3: Average empirical Cramér–von Mises distances (non-normalized) between the CDFs of the estimated complete data GEV and the proposed modelling approach computed over 100010001000 replications (“mle-obs” - observed likelihood, “mle-uncond” - unconditional weighted likelihood, “mle-cond” - conditional weighted likelihood”, “mom-obs” - observed moments, “mom-uncond” - unconditional weighted moments, “mom-cond” - conditional weighted moments) for MAR data generated from an t distribution (df=5) with blocks containing 505050 or 100100100 observations. Simulation parameters: “sims” - number of blocks, “apm” - average proportion of missingness across the series.